OpenStudy (anonymous):

Is (x−5) a factor of the polynomial p(x)=−x​^2​​−3x+10?

OpenStudy (zale101):

To simply know if it's a factor, try factoring the polynomial. Factor out -1 \(p(x)=-(x^2+3x-10)\) What two numbers when added you get +3 and when multiplied you get -10?

OpenStudy (anonymous):

im not sure what im supposed to do with this question. i dont understand :/

OpenStudy (zale101):

What are the factors of -10?

OpenStudy (anonymous):

1 2 5 10

OpenStudy (zale101):

Good. -5 * 2 =-10 but if -5+2 =-3 and we have +3 in our function \(p(x)=-(x^2\color{red}{+3}x-10)\)

OpenStudy (zale101):

If we hanged the negative signs, maybe we could get +3 so if 5*-2=-10 and if 5+(-2)=+3 Yes, so the factors of \(P(x)=-(x+5)(x-2)\) which can also be written like this if -1 is distributed to (x+5) \(P(x)=(-x-5)(x-2)\)

OpenStudy (zale101):

Do you think (x-5) is a factor?

OpenStudy (anonymous):

no

OpenStudy (zale101):

We have (x+5), it is not the same as (x-5) So (x-5) is not a factor.

OpenStudy (anonymous):

thank you

OpenStudy (zale101):

No problem